### v1

# Bridge Sampling Diagnostics

Abstract
In Bayesian statistics, the marginal likelihood is used for model selection and averaging, yet it is often challenging to compute accurately for complex models. Approaches such as bridge sampling, while effective, may suffer from issues of high variability of the estimates. We present how to estimate Monte Carlo standard error (MCSE) for bridge sampling, and how to diagnose the reliability of MCSE estimates using Pareto-$\hat{k}$ and block reshuffling diagnostics without the need to repeatedly re-run full posterior inference. We demonstrate the behavior with increasingly more difficult simulated posteriors and many real posteriors from the posteriordb database.
1 Introduction
The marginal likelihood is commonly used in Bayesian statistics. The purpose of this paper is not to advocate the use of marginal likelihood for model selection, but as marginal likelihood is used anyway, the goal is to help reducing the amount of computation used with practical diagnostics. Let $\mathcal{M}$ represent the Bayesian model of interest. The marginal likelihood given the observed data $\mathbf{y}$ is
$p(\mathbf{y}\mid\mathcal{M})=\int_{\Theta}p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})\,d\theta,$ (1)
where $p(\mathbf{y}\mid\theta,\mathcal{M})=\prod_{i=1}^{n}p(y_{i}\mid\theta,\mathcal{M})$ is the likelihood of $\theta$ given observed data $y_{i}$ and the model $\mathcal{M}$, $p(\theta\mid\mathcal{M})$ is the prior distribution of the parameters $\theta$, and $\Theta$ is the parameter space. For all but the simplest cases, this integral becomes intractable, requiring numerical methods to approximate it.
We examine the finite sample size accuracy of bridge sampling algorithm for estimating the marginal likelihood $p(\mathbf{y}\mid\mathcal{M})$. We (i) derive a closed-form expression for the Monte-Carlo standard error (MCSE) of the bridge sampling estimator, and (ii) introduce two complementary diagnostics–the Pareto-$\hat{k}$ tail index and a block-reshuffling bootstrap–to assess the adequacy of that MCSE estimate given finite MCMC sample size.
2 Bridge Sampling
One of the fastest and most reliable approaches to estimate the marginal likelihood is bridge sampling, first introduced by Bennett, (1976) and nicely reviewed by Gronau et al., (2017). The fundamental principle behind bridge sampling is the following identity
$1=\frac{\int p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})h(\theta)g(\theta)}{\int p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})h(\theta)g(\theta)},$ (2)
which allows $p(\mathbf{y}\mid\mathcal{M})$ to be written as
$p(\mathbf{y}\mid\mathcal{M})=\frac{\mathbb{E}_{g(\theta)}\left[p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})h(\theta)\right]}{\mathbb{E}_{\text{post}}\left[h(\theta)g(\theta)\right]},$ (3)
which consequently permits an approximation of the form
$\hat{p}(\mathbf{y}\mid\mathcal{M})=\frac{\frac{1}{S_{2}}\sum_{i=1}^{S_{2}}p(\mathbf{y}\mid\tilde{\theta}_{i},\mathcal{M})p(\tilde{\theta}_{i}\mid\mathcal{M})h(\tilde{\theta}_{i})}{\frac{1}{S_{1}}\sum_{j=1}^{S_{1}}h(\theta^{*}_{j})g(\theta^{*}_{j})},$ (4)
where $S_{1}$ is the number of Markov chain Monte Carlo draws $\theta^{*}_{j}$ from the posterior distribution $p(\theta\mid\mathbf{y},\mathcal{M})$, $S_{2}$ is the number of draws $\tilde{\theta}_{i}$ from a proposal distribution $g(\theta)$ with known normalization constant, and $h(\theta)$ is the bridge function.
To ensure the effectiveness of bridge sampling, $g(\theta)$ has to closely approximate $p(\theta\mid\mathbf{y},\mathcal{M})$, be easy to sample from, and have a known normalization constant. For these reasons, it is typically chosen to be a multivariate normal distribution, denoted by $\operatorname{N}(\mu,\Sigma)$, where the mean vector $\mu$ and covariance matrix $\Sigma$ are estimated using additional posterior draws. Usually the original posterior draws are divided in two sets, with one set used as $\theta^{*}_{j}$ and the other set used to estimate the parameters of the proposal distribution $g(\theta)$. Instead of the normal distribution, other distributions with known normalization constants can be used, or the posterior may be warped to resemble more closely a normal distribution, and the marginal likelihood is then computed based on this warped distribution, which preserves the same normalizing constant as the original distribution.
The bridge function $h$ typically used is the one proposed by Meng and Wong, (1996):
$h(\theta)\propto\frac{1}{s_{1}p(\mathbf{y}\mid\theta,\mathcal{M})p(\theta\mid\mathcal{M})+s_{2}p(\mathbf{y}\mid\mathcal{M})g(\theta)},$ (5)
where
$s_{1}\coloneqq\frac{S_{1}}{S_{1}+S_{2}}\quad\text{and}\quad s_{2}\coloneqq\frac{S_{2}}{S_{1}+S_{2}}.$
Because (5) depends on the marginal likelihood $p(\mathbf{y}\mid\mathcal{M})$, the bridge sampling estimator is refined through an iterative scheme:
$\hat{p}^{(t+1)}(\mathbf{y}\mid\mathcal{M})=\frac{\dfrac{1}{S_{2}}\sum_{i=1}^{S_{2}}\dfrac{l_{2,i}}{s_{1}l_{2,i}+s_{2}\hat{p}^{(t)}(\mathbf{y}\mid\mathcal{M})}}{\dfrac{1}{S_{1}}\sum_{j=1}^{S_{1}}\dfrac{1}{s_{1}l_{1,j}+s_{2}\hat{p}^{(t)}(\mathbf{y}\mid\mathcal{M})}},\qquad\text{with}\;\hat{p}^{(0)}(\mathbf{y}\mid\mathcal{M})=1.$ (6)
where
$l_{1,j}\coloneqq\frac{p(\mathbf{y}\mid\theta^{*}_{j},\mathcal{M})p(\theta^{*}_{j}\mid\mathcal{M})}{g(\theta^{*}_{j})},$
and
$l_{2,i}\coloneqq\frac{p(\mathbf{y}\mid\tilde{\theta}_{i},\mathcal{M})p(\tilde{\theta}_{i}\mid\mathcal{M})}{g(\tilde{\theta}_{i})}.$
The iteration process continues until the magnitude of the update falls below a predefined threshold and, to prevent the algorithm from oscillating indefinitely, it is also set to terminate after a specified maximum number of iterations.
Unlike plain importance sampling, bridge sampling is less sensitive to the mismatch in tail thickness in proposal and posterior distribution. Specifically, both ratios in the numerator and denominator in Equation (6) are bounded by construction (see, e.g., Gronau et al., , 2017). The bounded ratios imply that both the numerator and denominator have finite variance, and thus the whole estimator has finite variance.
3 Monte Carlo Standard Error
As the first step to assess the accuracy of the bridge sampling estimate, we propose to compute Monte Carlo standard error (MCSE). The estimator $\hat{p}(\mathbf{y})$ given in Equation (4) is a ratio of two sample means:
$\hat{p}(\mathbf{y}\mid\mathcal{M})=\frac{\bar{N}}{\bar{D}},$
with
$\bar{N}$ $=\frac{1}{S_{2}}\sum_{i=1}^{S_{2}}N_{i},\quad N_{i}\,\coloneqq\,p(\mathbf{y}\mid\tilde{\theta}_{i},\mathcal{M})\,p(\tilde{\theta}_{i}\mid\mathcal{M})\,h(\tilde{\theta}_{i}),$
$\bar{D}$ $=\frac{1}{S_{1}}\sum_{j=1}^{S_{1}}D_{j},\quad D_{j}\,\coloneqq\,h(\theta^{*}_{j})\,g(\theta^{*}_{j}).$
To compute the variance of $\hat{p}(\mathbf{y}\mid\mathcal{M})$, we use the delta method, thus, the variance estimate is given by:
$\operatorname{Var}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)\approx\frac{1}{S_{1}}\left(\frac{\bar{N}}{\bar{D}}\right)^{2}\left(\frac{\operatorname{Var}(\bar{N})}{\bar{N}^{2}}+\frac{\operatorname{Var}(\bar{D})}{\bar{D}^{2}}-2\,\frac{\operatorname{Cov}(\bar{N},\bar{D})}{\bar{N}\,\bar{D}}\right),$ (7)
assuming that $S_{1}=S_{2}$. As $\bar{N}$ and $\bar{D}$ are sample means, their variances and covariance are:
$\operatorname{Var}(\bar{N})$ $=\frac{\operatorname{Var}(N_{i})}{S_{2}},$
$\operatorname{Var}(\bar{D})$ $=\frac{\operatorname{Var}(D_{j})}{S_{1}},$
$\operatorname{Cov}(\bar{N},\bar{D})$ $=\frac{\operatorname{Cov}(N_{i},D_{j})}{S_{1}}.$
The draws $\tilde{\theta}_{i}$ and $\theta^{*}_{j}$ are independent, and thus $\operatorname{Cov}(\bar{N},\bar{D})=0$. Substituting back, we get
$\operatorname{Var}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)\approx\frac{1}{S_{1}}\left(\frac{\bar{N}}{\bar{D}}\right)^{2}\left(\frac{\operatorname{Var}(N_{i})}{S_{2}\,\bar{N}^{2}}+\frac{\operatorname{Var}(D_{j})}{S_{1}\,\bar{D}^{2}}\right).$ (8)
If $\theta^{*}_{j}$ are drawn using MCMC, terms $D_{j}$ may also have significant autocorrelation and estimate for $\operatorname{Var}(\bar{D})$ should then take that into account. In the experiments we used MCMC effective sample size estimate by Vehtari et al., (2021). Finally
$\operatorname{MCSE}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)=\sqrt{\operatorname{Var}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)}.$ (9)
To compute the MCSE of the logarithm of the marginal likelihood estimate, $\log\hat{p}(\mathbf{y}\mid\mathcal{M})$, we use the property of the log-normal distribution to get
$\operatorname{Var}\left(\log\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)=\log\left(1+\frac{\operatorname{Var}\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)}{\left(\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)^{2}}\right).$
Then
$\operatorname{MCSE}\left(\log\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)=\sqrt{\operatorname{Var}\left(\log\hat{p}\left(\mathbf{y}\mid\mathcal{M}\right)\right)}.$ (10)
The MCSE quantifies the uncertainty in the estimated (log) marginal likelihood due to the finite sample sizes $S_{1}$ and $S_{2}$, conditional on the proposal and the final bridge function when the iteration ends. In cases when the proposal is very different from the posterior or when the iterative algorithm fails to converge there can be additional variation. In Section 5, we present a block reshuffling approach which takes into account all the parts of the bridge sampling algorithm. However, based on the experiments reported in Section 6, the presented MCSE estimator alone is often sufficiently accurate.
4 Pareto-$\hat{k}$ Diagnostic
Let
$N_{i}\;=\;p(\mathbf{y}\mid\tilde{\theta}_{i},\mathcal{M})\,p(\tilde{\theta}_{i}\mid\mathcal{M})\,h(\tilde{\theta}_{i}),\qquad D_{j}\;=\;h(\theta^{*(j)})\,g(\theta^{*(j)}),$
so that the bridge estimator can be expressed as
$\hat{p}(\mathbf{y}\mid\mathcal{M})\;=\;\frac{\bar{N}}{\bar{D}},\qquad\bar{N}=\frac{1}{S_{2}}\sum_{i=1}^{S_{2}}N_{i},\qquad\bar{D}=\frac{1}{S_{1}}\sum_{j=1}^{S_{1}}D_{j}.$
Because $h(\cdot)$, $g(\cdot)$ and $p(\mathbf{y}\mid\cdot,\mathcal{M})\,p(\cdot\mid\mathcal{M})$ are all non-negative and bounded above on $\Theta$, there exist deterministic constants $B_{N},B_{D}<\infty$ such that
$0\;\leq\;N_{i}\leq B_{N},\qquad 0\;\leq\;D_{j}\leq B_{D}\quad\text{a.s.}$
Consequently, moments of distributions of $N_{i}$ and $D_{j}$ are all finite, and standard results (law of large numbers, delta method) guarantee that
$\sqrt{S_{1}}\!\left(\hat{p}(\mathbf{y}\mid\mathcal{M})-p(\mathbf{y}\mid\mathcal{M})\right)\;\xrightarrow{d}\;\operatorname{N}\bigl{(}0,\sigma^{2}\bigr{)},$
so, the variance-based Monte-Carlo standard error (MCSE) is asymptotically exact.
In practice the upper bounds $B_{N}$ and $B_{D}$ can be so large that, for realistic sample sizes $S_{1},S_{2}$, observations near either bound are exceedingly rare. The resulting finite sample behavior can be similar to that of heavy-tailed, or even infinite-variance, distributions: a single extreme realization may dominate $\bar{N}$ or $\bar{D}$ and inflate the estimator’s variance long before asymptotic behavior kicks in. Standard MCSE formulas, which assume finite variance, therefore give no early warning of this pre-asymptotic hazard.
[…]
$M\;=\;\Bigl{\lfloor}\min\{\,0.2\,S,\;3\sqrt{S}\}\Bigr{\rfloor},\qquad u=Z_{(M)},\qquad E_{i}=Z_{(i)}-u,\;i=1,\dots,M.$
The excesses $\{E_{i}\}$ are modeled with a generalized Pareto distribution
$p(x\mid u,\sigma,k)=\begin{cases}\dfrac{1}{\sigma}\left(1+k\,\dfrac{x}{\sigma}\right)^{-\left(\tfrac{1}{k}+1\right)},&k\neq 0,\;x\geq 0,\\[10.0pt] \dfrac{1}{\sigma}\exp\bigl{(}-x/\sigma\bigr{)},&k=0,\;x\geq 0,\end{cases}$
and the shape parameter $k$ is estimated using the fast approximate Bayesian computation proposed by Zhang and Stephens, (2009), yielding the scalar diagnostic $\hat{k}$. A value
$\hat{k}<0.5\quad\bigl{(}\text{or }0.7\text{ with smoothing}\bigr{)}$
indicates that the empirical mean $\bar{Z}$ behaves well in finite samples ($\hat{k}<0$ corresponding to a bounded distribution), whereas $\hat{k}$ above those thresholds signals behavior typical for heavy-tailed distributions. As the distributions are known to be bounded, increasing the sample sizes $S_{1}$ and $S_{2}$ will eventually lead to $\hat{k}<0$, although that may require impractically large sample sizes.
[…]
$\hat{k}_{N}\;=\;\hat{k}\bigl{(}\{N_{i}\}\bigr{)},\qquad\hat{k}_{D}\;=\;\hat{k}\bigl{(}\{D_{j}\}\bigr{)}.$
This procedure implicitly assumes tail-independence between $N_{i}$ and $D_{j}$, which is true conditionally on the final bridge function after the final iteration as the draws from the posterior and the proposal are independent.
Empirical results in Section 6 show that the component-wise $\hat{k}$ values are informative for bridge sampling and are thus a useful, but not exhaustive safeguard, as they reliably flag many heavy-tail pathologies in bridge sampling. Still, practical usefulness is limited as the tail shape can be far from Pareto shaped.
5 Block Reshuffling Diagnostic
The MCSE estimate discussed in Section 3 does not take into account potential extra variation due to two algorithm parts: 1) the original posterior draws are split in half with the second half used to estimate the parameters of the proposal distribution, and 2) the potential non-convergence of the iterative updates. The latter can happen if the proposal distribution happens to be bad and the empirical mean $\bar{N}$ or $\bar{D}$ is unstable.
To take into account the effect of these algorithm parts we propose to use a bootstrap style approach. Instead of regular bootstrap, the original posterior draws are divided in blocks and the blocks are shuffled before dividing the posterior draws in half and continuing with the algorithm as usual. The rationale behind using block reshuffling, as opposed to simply random shuffling of the draws, lies in the preservation of this autocorrelation structure, which influences the magnitude of Monte Carlo error. Re-shuffling and re-running bridge sampling $R>1$ times, we get a set of (log) marginal likelihood estimates $\{\hat{p}_{r}(\mathbf{y}\mid\mathcal{M})\}_{r=1}^{R}$. Standard deviation of these can be used as Monte Carlo standard error estimate.
Compared to actually rerunning MCMC, block re-shuffling is expected to be slightly optimistic as the reshuffled sets are overlapping and thus have dependency. Experimental results confirm usefulness of the block reshuffling and the slight optimistic bias.
The Pareto-$\hat{k}$ diagnostic can also be used to diagnose the tail shape of the distribution $\hat{p}_{r}(\mathbf{y}\mid\mathcal{M})$}. Based on the experiments in Section 6, Pareto-$\hat{k}$ for the block reshuffling estimates is more reliable than directly for the bridge sampling numerator and denominator, but has the additional cost of running bridge sampling several times.
[…]
6 Experimental Results
This section presents the empirical evaluation of the bridge sampling method, along with the proposed diagnostics. We begin by introducing a simple toy example, which serves to illustrate the conditions under which the bridge sampler performs effectively, as well as the circumstances in which it has too high variance to be useful. Following this, we present results using many real models and datasets sourced from the posterior database posteriordb, introduced by Magnusson et al., (2025).
6.1 Sparse Linear Regression with Increasing Number of Covariates
We begin by exploring the performance of bridge sampling within the context of linear regression with a highly non-normal posterior, and particularly focusing on how the method behaves as the number of covariates increases and the posterior gets higher dimensional and more non-normal.
The simulated data for this analysis were generated as follows. Given $n=100$ observations, the design matrix $\mathbf{X}$ was created with $k$ covariates, where $k$ varies from 10 to 103. Each element of $\mathbf{X}$ was drawn from a standard normal distribution and scaled by $\frac{1}{\sqrt{n\times k}}$:
$\mathbf{X}=\frac{1}{\sqrt{n\times k}}\mathbf{Z},\quad\text{where}\quad\mathbf{Z}\sim\mathcal{N}(0,1)^{n\times k}.$
The regression coefficients $\boldsymbol{\beta}$ were sampled from a standard normal distribution, with the last coefficient set to zero:
$\boldsymbol{\beta}=(\beta_{1},\beta_{2},\dots,\beta_{k-1},0)^{\top},\quad\text{where }\beta_{i}\sim\mathcal{N}(0,1).$
The response variable $\mathbf{y}$ was generated as a linear combination of the covariates with added normal noise:
$\mathbf{y}=\mathbf{X}\boldsymbol{\beta}+\frac{1}{\sqrt{n}}\boldsymbol{\epsilon},\quad\text{where }\boldsymbol{\epsilon}\sim\mathcal{\mathcal{missing}}{N}(0,2).$
The model is
$\mathbf{y}=\mathbf{X}\boldsymbol{\beta}+\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I}).$
To induce a non-normal posterior, a regularized horseshoe prior $\text{RHS}(\tau,\lambda_{i})$ (Piironen and Vehtari, , 2017) was used for the regression coefficients $\boldsymbol{\beta}$. The RHS prior was implemented as a scale mixture of normals:
$\beta_{i}\sim\mathcal{N}(0,\tau^{2}\lambda_{i}^{2}),\quad\lambda_{i}\sim\text{Cauchy}^{+}(0,1),\quad\tau\sim\text{Cauchy}^{+}(0,1).$
For the experiment, it doesn’t matter that the prior does not match how the data is generated. The important aspect is to make the posterior challenging for bridge sampling using a normal proposal distribution.
The posterior sampling was done using the rstanarm function stan_glm (Goodrich et al., , 2024). Bridge sampling was done using bridge_sampler function in the bridgesampling package (Gronau et al., , 2020). We sampled 4000 posterior draws from each posterior, which is the default in rstanarm.
Figure 1 shows the standard deviations of log marginal likelihood and marginal likelihood estimates, and the MCSE estimate for the first. As the number of covariates increases, the posterior gets more high-dimensional and non-normal, increasing the variability of the bridge sampling estimates of the (log) marginal likelihood. The MCSE estimate starts to fail around 50 covariates. Standard deviation estimation via block reshuffling slightly underestimates the standard deviation after about 60 covariates.
Figure 1: Top plot shows the standard deviation of the log marginal likelihood estimate (in log scale) from repeated running of $MC$ and bridge sampling (blue), the standard deviation of log marginal likelihood estimate from repeated running of bridge sampling with block reshuffling (green), and the MCSE estimate from a single bridge sampling estimate (red). The MCSE estimate starts to fail after about 45 covariates. The bottom plot shows the standard deviation (normalized by the median marginal likelihood given the number of covariates) of the marginal likelihood estimate (in log scale) from repeated running of $MC$ and bridge sampling (blue) and the standard deviation of the log marginal likelihood estimate from repeated running of bridge sampling with block reshuffling (green). Standard deviation estimation via block reshuffling slightly underestimates the standard deviation after about 60 covariates. The correspondence between the standard deviation for the log marginal likelihood and the standard deviation for the marginal likelihood is very close to the correspondence between the standard deviations of normal and log normal distributions.
To give some idea of the scale of the potential error, we can consider using the (log) marginal likelihood for model comparison. Based on the experiments, the uncertainty in the log marginal likelihood estimates is quite well approximated by a normal distribution. With bridge sampling, the (log) marginal likelihood estimates for two models are independent, and if the MCSE are similar, then MCSE for difference of log marginal likelihoods is $\sqrt{2}$ times as big. If the true difference would be 0, then a log marginal likelihood MCSE of about 0.5 would lead to 6% of the repeated estimates to be smaller than 1/3 and 6% to be bigger than 3, corresponding to moderate evidence for either model depending on chance. Thus, a desirable MCSE for the log marginal likelihood would be less than 0.2 if the difference is expected to be small (naturally, this depends on the specific use case).
If using Stan for the posterior inference, we recommend to start with the default number of posterior draws (4000), and then adjust based on the observed MCSE. For simple cases, MCSE with the default number of draws may already be sufficient [^1] At the time of writing, the bridgesampling package documentation states “The computation of marginal likelihoods based on bridge sampling requires a lot more posterior draws than usual. A good conservative rule of thump is perhaps 10-fold more draws.”. If the estimated MCSE is too high, more posterior draws can be obtained. As the variance of the estimator is finite, we may assume in many cases that MCSE halves by quadrupling the number of posterior draws and use that rule to estimate how many posterior draws are required for the desired accuracy. If the bound for either numerator or denominator in the bridge sampling equation is far, then even more draws may be required for the desired accuracy (Vehtari et al., , 2024).
In this example the biggest standard deviation for the log marginal likelihood is 2.5, which in comparing models with no true difference would lead to 10% of the repeated estimates to be smaller than 1/100 and 10% to be bigger than 100, corresponding to extreme evidence for either model depending on luck. To get a MCSE of 2.5 reduced to around 0.2 would require at least 156 times more posterior draws, that is, more than 600 000 posterior draws (and might require more if the pre-asymptotic convergence rate is lower due a long tail and very large bound). Alternate proposal distributions may also improve the performance.
We next look at the use of the Pareto-$\hat{k}$ as a further diagnostic. Figure 2 shows the average Pareto-$\hat{k}$ diagnostic using the default number of the tail values (135 given 2000 draws, continuous lines). Pareto-$\hat{k}$s for both numerator (yellow) and denominator (blue) are below 0.5 when the number of covariates is less than 21, and thus, in those cases, we could trust MCSE estimates.
Figure 2: Pareto-$\hat{k}$ diagnostic in the linear regression example with the increasing number of covariates for the distribution of terms in the numerator (yellow) and denominator (blue). The plot shows the average over 100 independent MCMC runs of estimated $\hat{k}$ using the default number of tail draws (135) and only 10 tail draws with continuous and dashed lines, respectively.
However, Figure 1 did show that the MCSE estimate was accurate up to about 45 covariates. In this case, the tails of the numerator and denominator distributions are far from Pareto shape, and Pareto-$\hat{k}$ with the default number of tail values gives too high values. Figure 2 also shows the average Pareto-$\hat{k}$ diagnostic using only 10 tail values (dashed lines). Now Pareto-$\hat{k}$s for both numerator and denominator are below 0.5 when the number of covariates is less than 42, which better matches the range when the MCSE estimate starts to fail. This shows that even when the distributions are bounded, with a finite sample the tail draws can resemble draws from a thick tail leading to an unreliable variance-based MCSE estimate. In this case, the average Pareto-$\hat{k}$ diagnostic using only 10 tail values matches the average behavior, but for a single case the Pareto-$\hat{k}$ using only 10 tail values has too big variability to be a practical diagnostic. We leave it for future research to develop an alternative tail shape diagnostic that would work well for such bounded distributions that arise in bridge sampling.
6.2 posteriordb Posteriors
To gain further insights into the performance of bridge sampling under different conditions, we proceed to an examination of 41 models and posteriors from the posteriordb database (Magnusson et al., , 2025) plus a multi-component Gaussian process model of birthday data (Zhang et al., , 2022). The posterior distributions selected for this study are the same as those analyzed by Zhang et al., (2022) plus several additional high dimensional posteriors that have been included in posteriordb more recently. They include a representative set of real life posteriors with varying dimensionality, non-normality, funnelness, and multi-modality. The list of posteriors, the number of posterior dimensions (in unconstrained space), and short description of the model is provided in Appendix A. For many models the Stan code used is dropping the constant terms, so the estimated marginal likelihoods are also missing that constant, but that does not affect the estimation variability. We sampled 4000 posterior draws from each posterior using CmdStanR with the default options except using the Pathfinder algorithm (Zhang et al., , 2022) to initialize MCMC for improved convergence speed and reliability. We did check that the MCMC sampling performed reasonably for all posteriors, but in some cases there are divergences which indicate the possibility of biased estimates, which can also increase the bias and variability in the (log) marginal likelihood estimates. We did not compute true marginal likelihood for any of the posteriors, and it would be impossible for most of the posteriors, but compare the performance to the average over 100 independent repeated runs.
Figure 3 shows MCSE and standard deviation of log marginal likelihood estimates from block reshuffling vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. MCSE estimates are quite accurate for most of the posteriors while block reshuffling standard deviation tends to be slightly smaller than the MCMC standard deviation when the true standard deviation is also small. For 36/42 posteriors MCSE and MCMC standard deviation are below 0.2. For the Birthdays posterior we can estimate that 13 times more posterior draws, that is, at least 52 000 posterior draws would be required to get a MCSE of 0.2 or smaller. For the Presidential votes hierarchical Gaussian process posterior we can estimate that 500 times more posterior draws, that is, at least 2 million posterior draws would be required to get MCSE of 0.2 or smaller, which starts to be impractically big sample size.
Figure 3: posteriordb posteriors + Birthdays: The left plot shows estimated MCSE vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. The right plot shows standard deviation of log marginal likelihood estimates from block reshuffling vs standard deviation of log marginal likelihood estimates from 100 repeated MCMC runs. MCSE estimate has small bias except for a few posteriors with the highest variability. Block reshuffling underestimates the small variability estimates, but has small bias in case of high variability.
Figure 4 shows the strong dependency between the number of posterior dimensions and standard deviation of log marginal likelihood estimates. The increasing number of dimensions increases the variability, but only posteriors with more than 300 dimensions have standard deviation larger than 0.2, which can be considered to be an excellent result. The variation within the posteriors having the same dimensionality can be explained by a) how different the posterior is from normal, and b) how efficient MCMC sampling is. For example, eight schools with centered parameterization is known to have a strong funnel-shaped posterior which Stan’s current MCMC algorithm is not able to sample efficiently.
Figure 4: posteriordb posteriors + Birthdays: The number of posterior dimensions vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. The estimate variability tends to increase with the number of dimensions, which is natural due to the curse of dimensionality. In addition there is variation depending on how non-normal the posterior distribution is.
7 Discussion
We have proposed using the usual Monte Carlo standard error estimate to diagnose the reliability of (log) marginal likelihood estimates obtained by bridge sampling. We have demonstrated that this simple MCSE estimate is sufficiently accurate for practical purposes. When a MCSE estimate fails, it is likely that the provided MCSE estimate is anyway so big that it is indicating the need to run more MCMC iterations. In the case of high MCSE estimates, further accuracy can be obtained by running block reshuffling. The error in the log marginal likelihood seems to have a close to normal distribution, and the error in the marginal likelihood seems to have a close to log-normal distribution. Although the log-normal distribution has all finite moments, with finite sample size, the behavior can be indistinguishable from a distribution with infinite variance. The Pareto-$\hat{k}$ diagnostic can provide useful information about the reliability of MCSE estimates for bridge sampling, but its practical benefit is small due to the Pareto tail assumption not being sufficiently accurate for the relevant ratio distributions. We leave it for future research to investigate other tail shape measures that would work better with tail shapes arising in bridge sampling and could provide a more accurate diagnostic for how many more iterations would be needed for a desired accuracy of the (log) marginal likelihood estimate.
Acknowledgements
We thank Ozan Adıgüzel for making some of the initial experiments (not shown in this paper), and Noa Kallioinen and Anna Riha for helpful comments and feedback. We acknowledge the support by the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence, Research Council of Finland project (340721), and The Aalto Science Institute international summer research internship programme, as well as the computational resources provided by the Aalto Science-IT project.
Appendix A posteriordb Posteriors and Birthdays Posterior
We used 41 posteriors from posteriordb and a multi-component Gaussian process model (gpbf6), which was used by Zhang et al., (2022). Table A lists the posteriordb IDs, the posterior names used in the Figures in this paper, the number of posterior dimensions (in unconstrained space), and the short description of the model. The code for running the posterior inference and bridge sampling for all these and also for the simulated toy example is available at https://github.com/GiorgioMB/bridgesampling_paper_code.
posteriordb id plot names D description
[…]
arma-arma11 ARMA(1,1) 4 ARMA
bball_drive_event_1-hmm_drive_1 Bball HMM 6 Hidden Markov model
bball_drive_event_0-hmm_drive_0 Bball0 HMM 6 Hidden Markov model
- Birthdays GP 438 Multi-component GP time series model (gpbf6)
butterfly-multi_occupancy Butterfly 106 Multiple Species-site occupancy model
diamonds-diamonds Diamonds 26 Multiple highly correlated predictors log-log model
dogs-dogs Dogs 3 Logistic mixed effects model
dogs-dogs_log Dogs Log 2 Logarithmic mixed effects model
earnings-logearn_interaction Earnings 5 Multiple predictors interacting log-linear model
eight_schools-eight_schools_centered 8-Schools CP 10 A centered hiearchical model for 8 schools
eight_schools-eight_schools_noncentered 8-Schools NCP 10 A non-centered hiearchical model for 8 schools
garch-garch11 GARCH 4 Generalized autoregressive conditional heteroscedastic model
GLM_Binomial_data-GLM_Binomial_model GLM Binomial 3 Success rate of peregrine broods
GLM_Poisson_Data-GLM_Poisson_model GLM Poisson 4 Poisson GLM for modeling a population of peregrines
GLMM_data-GLMM1_model GLMM Poisson 1 237 GLMM for peregrine population size
GLMM_Poisson_data-GLMM_Poisson_model GLMM Poisson 2 45 GLMM for peregrine population size with random site and year effects
gp_pois_regr-gp_regr GP Normal 3 Gaussian process regression
gp_pois_regr-gp_pois_regr GP Poisson 13 Gaussian process Poisson regression
hmm_example-hmm_example HMM 4 Hidden Markov model
hudson_lynx_hare-lotka_volterra Hudson Lynx Hare 8 Lotka-Volterra Model
irt_2pl-irt_2pl 2PL IRT 144 Two-parameter logistic item response theory model
fims_Aus_Jpn_irt-2pl_latent_reg_irt Japan IRT 531 Two-parameter logistic item response theory model
low_dim_gauss_mix-low_dim_gauss_mix Low Dim Gauss Mix 5 A Two-dimensional Gaussian mixture model
lsat_data-lsat_model LSAT Rasch 1006 Random effects (Rasch) model for true difficulty of LSAT questions
Mb_data-Mb_model Capture-recapture Mb 3 Inferring population size considering immediate trap-response
mcycle_gp-accel_gp Mcycle GP 66 Heteroscedastic Gaussian processes
Mh_data-Mh_model Capture-recapture Mh 388 Logistic-normal heterogeneity model
Mtbh_data-Mtbh_model Capture-recapture Mtbh 154 Capture-recapture model where detection probability varies by occasion
Mth_data-Mth_model Capture-recapture Mth 394 Capture-recapture model where detection probability varies by occasion and heterogeneity
nes2000-nes NES2000 10 Multiple predictor linear model
one_comp_mm_elim_abs-one_comp_mm_elim_abs PKPD One Comp 4 One compartment PKPD model
radon_all-
radon_hierarchical_intercept_noncentered Radon Hierarchical 391 A county intercept and county level covariate for Radon data (non-centered)
radon_all-
radon_variable_intercept_slope_noncentered Radon Intercept 777 Variable intercept and slope hiearchical for Radon data (noncentered)
radon_all-
radon_partially_pooled_noncentered Radon Pooled 389 Hiearchical intercept model for Radon data (non-centered)
radon_all-
radon_variable_slope_noncentered Radon Slope 390 Variable slope hiearchical for Radon data (non-centered)
sat-hier_2pl SAT Rasch 670 Hierarchical two-parameter logistic item response model
sblrc-blr sblrc 6 A Bayesian linear regression model with vague priors
science_irt-grsm_latent_reg_irt Science IRT 408 Rating scale and generalized rating scale models with latent regression
state_wide_presidential_votes-hierarchical_gp Pres-votes hier-GP 933 Hierarchical Gaussian process
Survey_data-Survey_model Survey 1 Inferring the return rate and number of surveys from observed returns
timssAusTwn_irt-gpcm_latent_reg_irt timssAusTwn IRT 530 Partial credit and generalized partial credit models with latent regression